# Attention Surgery: An Efficient Recipe to Linearize Your Video Diffusion Transformer

### v2

@@ 15 Linear
Comparison with SOTA. Tab. 1 shows a comparison of one of the variations of our model with the state-of-the-art methods on VBench. Results indicate that our hybrid operated models are competitive with state-of-the-art efficient video diffusion models, while one of the variations (15R2) is equivalent to the original Wan2.1* model it’s based on. On VBench-2.0 15R2 gets total score of 55.1, comparable to the original 56.0. Please refer to Supplementary for full set of evaluations.
@@ A.1 Training Details and Hyperparameters
Pretraining distillation stage. We train each block independently and all the parameters are frozen except for the and , for which we use the AdamW optimizer Loshchilov & Hutter (2019), batch size of 1 and a learning rate of 1e-3, with the value distillation objective, as detailed in equation 7 to train. To extract teacher activations for distillation, we sample using 50 denoising steps with a guidance scale of 5, employing the Euler Ancestral Discrete Scheduler to integrate the reverse diffusion process.
@@ A.2 Qualitative samples
Figures 8 to 25 show uniformly spaced frames from videos generated by the original Wan2.1 1.3B and different variations of our hybrid attention models (15R2, 15R4, 15R8, 20R4, and 20R8), for 18 different prompts on the original 480832 resolution. All the videos corresponding the demonstrated frames, are available as video files in the attached supplementary materials.
A.3 Detailed VBench Comparison
@@
A.4 Detailed VBench-2.0 Comparison
Tabs. 9, 9 and 9 demonstrate fine-grained results on the recent VBench-2.0 benchmark at original resolution of 480832. We generated videos with the original Wan2.1 1.3B model and our 15R2 modification using the same sampler hyperparameters. We observe that our hybrid model experiences an insignificant drop in performance as measured by Total score.

### v3

@@ 15 Linear
Comparison with SOTA. Tab. 1 shows a comparison of one of the variations of our model with the state-of-the-art methods on VBench. Results indicate that our hybrid operated models are competitive with state-of-the-art efficient video diffusion models, while one of the variations (15R2) is equivalent to the original Wan2.1* model it’s based on. As summary of VBench-2.0 comparisons are presented in Tab 4. Please refer to Supplementary for the full set of evaluation dimensions.
| Model | VBench-2.0 | |||||
| Total | Hum.Fid. | Creativity | Control. | Com.sense | Physics | |
| Wan2.1 1.3B | 56.0 | 80.7 | 48.7 | 34.0 | 63.4 | 53.8 |
| CogVideoX-1.5 5B | 53.4 | 72.1 | 43.7 | 29.6 | 63.2 | 48.2 |
[…]
@@
Computational Efficiency. Fig. 2 right, shows the dramatic savings in compute burden in terms of FLOPs and on-mobile latencies, as we increase the video duration on 320480 frame resolution. For instance, this results in 6 faster inference per block for videos of 7.5s. Clearly, higher resolutions such as 480p or 720p only increase the gap as they bring the comparison to a regime with larger number of tokens. Memory read/write comparisons are available in the appendix.
@@ A.1 Training Details and Hyperparameters
Pretraining distillation stage. We train each block independently and all the parameters are frozen except for the and , for which we use the AdamW optimizer loshchilov2019decoupled, batch size of 1 and a learning rate of 1e-3, with the value distillation objective, as detailed in equation 7 to train. To extract teacher activations for distillation, we sample using 50 denoising steps with a guidance scale of 5, employing the Euler Ancestral Discrete Scheduler to integrate the reverse diffusion process.
@@ A.2 Qualitative samples
Figures 9 to 26 show uniformly spaced frames from videos generated by the original Wan2.1 1.3B and different variations of our hybrid attention models (15R2, 15R4, 15R8, 20R4, and 20R8), for 18 different prompts on the original 480832 resolution. All the videos corresponding the demonstrated frames, are available as video files in the attached supplementary materials.
A.3 Memory Read/Write on Mobile
Table 8 shows the the memory read/write values for one DiT block in various methods in GB, as measured with the QNN runtime on a Snapdragon8-Gen4 SoC.
| Number of frames - Memory Read/Write (GB) | ||||||||
| 81 | 101 | 121 | 141 | |||||
| Attention Block | W | R | W | R | W | R | W | R |
| Softmax Flash Attention | 5.1 | 6.0 | 12.9 | 16.4 | 22.7 | 53.6 | OOM | OOM |
| HedgeHog Linear Attention | 5.7 | 8.1 | 7.0 | 10.1 | 6.9 | 11.3 | 8.0 | 13.2 |
| Attention Surgery - R8 | 6.3 | 10.1 | 5.2 | 10.9 | 6.4 | 13.2 | 7.8 | 35.2 |
A.4 Detailed VBench Comparison
@@
A.5 Detailed VBench-2.0 Comparison
Tabs. 11, 11 and 11 demonstrate fine-grained results on the recent VBench-2.0 benchmark at original resolution of 480832. We generated videos with the original Wan2.1 1.3B model and our 15R2 modification using the same sampler hyperparameters. We observe that our hybrid model experiences an insignificant drop in performance as measured by Total score.